OpenStudy (anonymous):

Given a=i+j, b=i-j, and u=2i+3j, express u in terms of a and b. [Hint: let u=ma+nb]

OpenStudy (ajprincess):

Let u=ma+nb 2i+3j=m(i+j)+n(i-j) =mi+mj+ni-nj =(m+n)i+(m-n)j Nw let us compare the coefficients of i and j m+n=2-(1) m-n=3-(2) (1)+(2)=> m+n+m-n=2+3 2m=5 From this m can be found. then plug in that value in either (1) or (2) and find n.

OpenStudy (anonymous):

Alright thankyou so much!

OpenStudy (ajprincess):

welcome:)

OpenStudy (anonymous):

wait i dont understand the part where you did m+n=2-(1) where did u get the 1

OpenStudy (ajprincess):

I just name the equation as (1)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

ohh i get it. Thankss :)

OpenStudy (ajprincess):

Welcome:)